OpenStudy (anonymous):

Nathan's brother says when he multiples a number by 6 and adds 5 to the result he gets 23. To figure out the number, what should Nathan do first?

OpenStudy (anonymous):

it is 3 u subtract 5 from 18 and then div. 6 from 18 and u get 5 5-23=18/6=3